# Enantiosensitive molecular compass

### v6

@@ 1 Introduction
Chiral molecular interactions are a remarkable example of a geometrically robust response active in living matter, maintained in complex and noisy environments and operating at ambient conditions [CISS, aiello2022chirality]. It has potential for applications, e.g., in quantum technologies [gaita2019molecular, atzori2019second, wasielewski2020exploiting, carretta2021perspective, heinrich2021quantum], with CISS being an example of useful functionality [chiesa2023chirality]. Originally observed as enantio-sensitive electron spin polarization upon transport through chiral biomolecules [ray1999asymmetric, ray2006chirality], CISS now encompasses a broad range of electronic processes where molecular chirality governs spin polarization [CISS, chiesa2023chirality]. Beyond molecular systems, solid-state materials with chiral crystal lattices exhibit a related but more complex influence on spin: they stabilize noncollinear spin textures such as skyrmion crystals - topologically nontrivial spin configurations whose formation and control are governed by the chirality of the lattice [muhlbauer2009skyrmion, nagaosa2013topological].
Many CISS-related phenomena are induced by light, suggesting the possibility of ultrafast control over spin-sensitive chemical processes via photoexcitation or photoionization. The interplay between spin and chirality naturally suggests spin-orbit coupling (SOC) as the underlying mechanism for spin selectivity. However, several experiments have shown that SOC alone cannot explain the observed effect [kettner2018chirality, mishra2013spin, kettner2016silicon]. Numerous theoretical approaches including scattering theory [yeganeh2009chiral, medina2012chiral, eremko2013spin, varela2013inelastic], tight-binding models [gutierrez2012spin, guo2012spin, matityahu2016spin, varela2016effective, peter_yelin_2024], density functional theory [maslyuk2018enhanced, diaz2018thermal, zollner2020insight], scattering off magnetic impurities [ghazaryan2020filtering], electron correlation models [fransson2019chirality], electron-phonon coupling [fransson2020vibrational, vittmann2023spin], non-adiabatic coupling [bian2021modeling, teh2022spin], and field-theoretical treatments [shitade2020geometric] have been proposed, yet a complete understanding remains elusive [CISS]. Diverse challenges for interpretation emerge due to the presence of multiple factors that potentially affect this process in experimental measurements thereby complicating its analysis, e.g., leads or substrates for molecules, and impurities, defects, or spurious fields for solids.
@@ 3 Quantifying spin-orientation locking
To quantify spin-chirality coupling in photoionization, we construct spin-resolved chiral electronic states in an Ar atom, rigorously incorporating electronic chirality and spin-orbit coupling. Inspired by analogous chiral hydrogenic states [ordonez2019propensity], these states combine excited-state orbitals:
@@
Figure 4: Enantio-sensitive spin-orientation locking under isotropic illumination of randomly oriented electronic states. (a) The Bloch pseudovector $\boldsymbol{\vec{S}}^{M}$ (internal directional bias) in the molecular frame for the chiral argon states. $\boldsymbol{\vec{S}}^{M}$ changes its direction in space as a function of the photoelectron momentum $k$. Trajectories traced by this vector are shown for $0<k<0.8\text{ Bohr}^{-1}$. (b) Degree of orientation for chiral states with $m=1$, $\mu=\pm\frac{1}{2}$ (violet and orange correspondingly), and averaged over spin orientation in initial state (black). (c) Degree of orientation for chiral states with m=1m=-1, $\mu=\pm\frac{1}{2}$ (red and green correspondingly) and averaged (black). The rapidly oscillating behavior at higher values of $k$ are due to the Fano resonances, leading up to the ionization threshold for the 3s electrons [Samson2002, Carlstroem2024spinpolspectral].
@@ 4 Conclusions and Outlook
Quantum correlations between the photoelectron spin and the molecular skeleton orientation are unique to chiral molecules. Importantly, they survive averaging over random molecular orientations, with the coherence between the spin-up and spin-down final components of the photoelectron recorded in the Bloch pseudovector resulting from the reduced density matrix. These correlations underlie the main microscopic mechanism of chirality induced spin selectivity. We identify the photoionization Bloch vector as a chiral molecular compass mediating spin-chirality coupling and offering significant control over the orientation of enantiomers in space. Conversely, the chiral molecular compass also gives rise to enantio-sensitive spin dynamics in photoexcited oriented molecules. Our approach highlights that both the direction and magnitude of the chiral molecular compass depend sensitively on photon energy, leading to either strong or weak spin polarization within the same chiral medium. We expect that the enantio-sensitive molecular compass is also relevant for understanding electron spin polarization in scattering from gas-phase chiral molecules [campbell1987electron, mayer1995experimental] - the phenomenon that predated the formulation of the CISS effect.
@@ Methods
Quantification of spin-chirality coupling in photoionization requires a system where both electronic chirality and spin-orbit coupling are rigorously accounted. To this end, we construct chiral electronic densities in an Ar atom corresponding to excitation into a chiral superposition of excited states, resolved on the excited electron spin. The eigenstates and associated spin-resolved photoionization dipole matrix elements of Ar atom were calculated using an atomic configuration-interaction singles treatment [carlstrom2022general1, carlstrom2022general2, Carlstroem2024spinpolspectral]:
@@
Chiral spin resolved electronic states in Ar are inspired by a similar spinless chiral superposition in the hydrogen atom [ordonez2019propensity], and in the case of hydrogenic wave-functions could be represented as:
@@
The vectors that appear on the right-hand side of Eq. (45) can now be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin expectation value $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [andrews1977three] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants.
@@
An estimate of the number of ‘head’ $N_{+\boldsymbol{\hat{s}}}$ and ‘tail’ $N_{-\boldsymbol{\hat{s}}}$ molecules corresponding to the averaged angle $\langle\boldsymbol{\hat{e}}^{L}\cdot\boldsymbol{\hat{s}}^{L}\rangle=\langle\cos\beta\rangle$ between the vector $\boldsymbol{\hat{e}}^{L}$ and “target” orientation axis $\boldsymbol{\hat{s}}^{L}$, can be performed by adopting the methods of Ref. [ordonez2023geometric] to model the angular distribution of oriented molecules. Let
@@ Appendix A Chiral superpositions of Argon excited states
In a close-coupling description of the excited states, the active electron is entangled with the ion. In the particle–hole basis employed in configuration-interaction singles [carlstrom2022general1, carlstrom2022general2], the excited states have contributions from multiple ion channels:
@@ A.1 State Factorization by Block Diagonalization in Orthogonal Subspaces
which we recognize as an eigenvalue problem. The minimization problem (62) may be solved using any standard non-linear solver; its variation (63) can be used to improve convergence. In our implementation, however, we employ Riemannian manifold optimization [Mogensen2018optim] of the minimization problem (61) as-is; the matrix $U$ is required to stay on the Stiefel manifold, i.e. the matrix manifold of matrices with mutually orthonormal columns. In Fig. 7, the optimized $U$ is shown, together with the transformed overlap matrix $UB$.
Figure 7: The left panel shows the optimized unitary transform matrix $U$. The right panel shows the overlap matrix $UB$ between rotated excited states and the spin-configurations, again grouped by the excited spin-orbitals.
@@ Appendix B Quantifying spin-orientation locking
The vectors that appear on the right-hand side of Eq. (77) can now be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin expectation value $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [andrews1977three] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants. For our purposes, the following vector identities will be relevant:

### v7

@@ 1 Introduction
Chiral molecular interactions are a remarkable example of a geometrically robust response active in living matter, maintained in complex and noisy environments and operating at ambient conditions [13, 1]. It has potential for applications, e.g., in quantum technologies [16, 3, 45, 9, 20], with CISS being an example of useful functionality [10]. Originally observed as enantio-sensitive electron spin polarization upon transport through chiral biomolecules [36, 37], CISS now encompasses a broad range of electronic processes where molecular chirality governs spin polarization [13, 10]. Beyond molecular systems, solid-state materials with chiral crystal lattices exhibit a related but more complex influence on spin: they stabilize noncollinear spin textures such as skyrmion crystals - topologically nontrivial spin configurations whose formation and control are governed by the chirality of the lattice [29, 30].
Many CISS-related phenomena are induced by light, suggesting the possibility of ultrafast control over spin-sensitive chemical processes via photoexcitation or photoionization. The interplay between spin and chirality naturally suggests spin-orbit coupling (SOC) as the underlying mechanism for spin selectivity. However, several experiments have shown that SOC alone cannot explain the observed effect [22, 27, 21]. Numerous theoretical approaches including scattering theory [46, 26, 12, 42], tight-binding models [19, 18, 24, 43, 35], density functional theory [23, 11, 48], scattering off magnetic impurities [17], electron correlation models [14], electron-phonon coupling [15, 44], non-adiabatic coupling [4, 41], and field-theoretical treatments [40] have been proposed, yet a complete understanding remains elusive [13]. Diverse challenges for interpretation emerge due to the presence of multiple factors that potentially affect this process in experimental measurements thereby complicating its analysis, e.g., leads or substrates for molecules, and impurities, defects, or spurious fields for solids.
@@ 3 Quantifying spin-orientation locking
To quantify spin-chirality coupling in photoionization, we construct spin-resolved chiral electronic states in an Ar atom, rigorously incorporating electronic chirality and spin-orbit coupling. Inspired by analogous chiral hydrogenic states [33], these states combine excited-state orbitals:
@@
Figure 4: Enantio-sensitive spin-orientation locking under isotropic illumination of randomly oriented electronic states. (a) The Bloch pseudovector $\boldsymbol{\vec{S}}^{M}$ (internal directional bias) in the molecular frame for the chiral argon states. $\boldsymbol{\vec{S}}^{M}$ changes its direction in space as a function of the photoelectron momentum $k$. Trajectories traced by this vector are shown for $0<k<0.8\text{ Bohr}^{-1}$. (b) Degree of orientation for chiral states with $m=1$, $\mu=\pm\frac{1}{2}$ (violet and orange correspondingly), and averaged over spin orientation in initial state (black). (c) Degree of orientation for chiral states with m=1m=-1, $\mu=\pm\frac{1}{2}$ (red and green correspondingly) and averaged (black). The rapidly oscillating behavior at higher values of $k$ are due to the Fano resonances, leading up to the ionization threshold for the 3s electrons [39, 8].
@@ 4 Conclusions and Outlook
Quantum correlations between the photoelectron spin and the molecular skeleton orientation are unique to chiral molecules. Importantly, they survive averaging over random molecular orientations, with the coherence between the spin-up and spin-down final components of the photoelectron recorded in the Bloch pseudovector resulting from the reduced density matrix. These correlations underlie the main microscopic mechanism of chirality induced spin selectivity. We identify the photoionization Bloch vector as a chiral molecular compass mediating spin-chirality coupling and offering significant control over the orientation of enantiomers in space. Conversely, the chiral molecular compass also gives rise to enantio-sensitive spin dynamics in photoexcited oriented molecules. Our approach highlights that both the direction and magnitude of the chiral molecular compass depend sensitively on photon energy, leading to either strong or weak spin polarization within the same chiral medium. We expect that the enantio-sensitive molecular compass is also relevant for understanding electron spin polarization in scattering from gas-phase chiral molecules [5, 25] - the phenomenon that predated the formulation of the CISS effect.
@@ Methods
Quantification of spin-chirality coupling in photoionization requires a system where both electronic chirality and spin-orbit coupling are rigorously accounted. To this end, we construct chiral electronic densities in an Ar atom corresponding to excitation into a chiral superposition of excited states, resolved on the excited electron spin. The eigenstates and associated spin-resolved photoionization dipole matrix elements of Ar atom were calculated using an atomic configuration-interaction singles treatment [7, 6, 8]:
@@
Chiral spin resolved electronic states in Ar are inspired by a similar spinless chiral superposition in the hydrogen atom [33], and in the case of hydrogenic wave-functions could be represented as:
@@
The vectors that appear on the right-hand side of Eq. (45) can now be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin expectation value $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [2] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants.
@@
An estimate of the number of ‘head’ $N_{+\boldsymbol{\hat{s}}}$ and ‘tail’ $N_{-\boldsymbol{\hat{s}}}$ molecules corresponding to the averaged angle $\langle\boldsymbol{\hat{e}}^{L}\cdot\boldsymbol{\hat{s}}^{L}\rangle=\langle\cos\beta\rangle$ between the vector $\boldsymbol{\hat{e}}^{L}$ and “target” orientation axis $\boldsymbol{\hat{s}}^{L}$, can be performed by adopting the methods of Ref. [32] to model the angular distribution of oriented molecules. Let
@@ Appendix A Chiral superpositions of Argon excited states
In a close-coupling description of the excited states, the active electron is entangled with the ion. In the particle–hole basis employed in configuration-interaction singles [7, 6], the excited states have contributions from multiple ion channels:
@@ A.1 State Factorization by Block Diagonalization in Orthogonal Subspaces
which we recognize as an eigenvalue problem. The minimization problem (62) may be solved using any standard non-linear solver; its variation (63) can be used to improve convergence. In our implementation, however, we employ Riemannian manifold optimization [28] of the minimization problem (61) as-is; the matrix $U$ is required to stay on the Stiefel manifold, i.e. the matrix manifold of matrices with mutually orthonormal columns. In Fig. 7, the optimized $U$ is shown, together with the transformed overlap matrix $UB$.
Figure 7: The left panel shows the optimized unitary transform matrix $U$. The right panel shows the overlap matrix UBUB between rotated excited states and the spin-configurations, again grouped by the excited spin-orbitals.
@@ Appendix B Quantifying spin-orientation locking
The vectors that appear on the right-hand side of Eq. (77) can now be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin expectation value $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [2] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants. For our purposes, the following vector identities will be relevant: